OpenStudy (anonymous):

The Millers plan on retiring soon. When they were first married they purchased 1,000 shares in a mutual fund that was at that time paying $10 a share. Now, thirty years later is is paying $78 a share. What is the percent increase for the fund? A) 6.8% B) 68% C) 680% D) 6,800%

OpenStudy (anonymous):

First we subtract 78 - 10

OpenStudy (anonymous):

I think it is C 680% I just tried to work it out

OpenStudy (anonymous):

Yeah, I got the same thing..

OpenStudy (anonymous):

78 - 10 = 68 68 / 10 = 6.8 6.8 * 100 = 680%